OpenStudy (anonymous):

Help with Radical Equations , Can you show work please? (3x)/(2x+1)-(4)/(2x-1)=1

OpenStudy (anonymous):

\[3x(2x-1)-4(2x+1)=(2x-1)(2x+1)\] \[6x^2-3x-8x-4=4x^2-1\] \[2x^2-11x-3=0\] \[x=(1/4)(11\pm \sqrt{145})\]

OpenStudy (anonymous):

Also, using the equation emileokada got: \[2x ^{2}-11x-3=0\] you can use the quadratic formula and get a better x. Quadratic Formula: \[x=-b \pm \sqrt{b ^{2}-4ac}\] all over 2a.